OpenStudy (anonymous):

which of the equations is represented in the following graph? 4x=y=10 3x+2y=12 4x+y=10 2x-3y=12 4x-y=10 2x+3y=12 x+4y=10 3x+2y=12

OpenStudy (anonymous):

here is the actual graph, which probably useful. lol

OpenStudy (rsadhvika):

notice that one line is falling, one line is rising

OpenStudy (rsadhvika):

falling line :- y-intercept = 4 slope = -4/6 = -2/3 so the equation wud be : y = -2/3x+4 3y=-2x+12 2x+3y = 12 --------------------------(1)

OpenStudy (rsadhvika):

check ur options, u have only one options with above equation.

OpenStudy (anonymous):

that makes sense! thanks!!!

OpenStudy (rsadhvika):

yw